OpenStudy (lilday12):

Evaluate the expression below. ( 1 + | -17 + 3 | ) ( -3 )

OpenStudy (jameshorton):

what yu think your answer is

OpenStudy (anonymous):

so it is saying 1+17+3 (3)

OpenStudy (lilday12):

i got 6.3

OpenStudy (anonymous):

\[(1+17+3)(3)\]

OpenStudy (anonymous):

so its 1+14 times 3

OpenStudy (anonymous):

\[( 1 + | -17 + 3 | ) \] \[-17+3=14 abs is always positive\] \[1+14 =15 \times -3 =-45\]

OpenStudy (texaschic101):

yep

OpenStudy (anonymous):

is this right @texaschic101

OpenStudy (texaschic101):

it is correct

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

this is fairly simple math

OpenStudy (texaschic101):

sometimes it can be tricky

OpenStudy (anonymous):

yea ik

OpenStudy (anonymous):

lol @texaschic101 i cant get the titles because my profile page dosnt load up on mycomputer or the phone